OpenStudy (anonymous):

Absolute value equation z-10=|7z|-16

OpenStudy (michele_laino):

we have to consider two cases, namely: \[\Large \begin{gathered} \left\{ \begin{gathered} z - 10 = 7z - 16 \hfill \\ z \geqslant 0 \hfill \\ \end{gathered} \right. \hfill \\ \hfill \\ \left\{ \begin{gathered} z - 10 = - 7z - 16 \hfill \\ z < 0 \hfill \\ \end{gathered} \right. \hfill \\ \end{gathered} \]